Arlene earns $170 per week less than Karen. The sum of their earnings is $1210 per week. How much does Arlene earn per week?
Answer by JulietG(1812)

Let A represent Arlene
Let K represent Karen
What do we know from the problem?
A + K = 1210
K = A + 170
Substitute the value from the second equation into the first.
A + K = 1210
A + (A+170) = 1210
2A + 170 = 1210
2A = 1040
A = 520
If A is 520, then K is 1210-520 or 690. 690 is 170 more than 520.
Karen makes $690/week.
Arlene makes $520/week.